OpenStudy (zion):

tan^2-cot^2+csc^2=? solutions?

OpenStudy (anonymous):

sec^2

OpenStudy (anonymous):

since csc^2-cot^2 = 1 & 1+ tan^2 is sec^2

OpenStudy (zion):

\[\tan ^{2}\theta-\cot^{2}\theta+\csc^{2}\theta=?\]

OpenStudy (anonymous):

cosec and sec have relations isnt it??..hw is cosec and cot hav??

OpenStudy (anonymous):

yeah.. csc2θ−cot2θ=1

OpenStudy (anonymous):

1+tan2θ=sec^2θ

OpenStudy (anonymous):

so your answer is sec^2 θ

OpenStudy (zion):

how did u type that theta symbol manoranjan?

OpenStudy (anonymous):

\theta